OpenStudy (anonymous):

What is the solution set of |x + 3| > 7? x < –10 or x > 4 x < 10 or x > –4 x > –10 or x < 4 x > 10 or x < –4

OpenStudy (anonymous):

Solve the following system of equations. 4x – 3y + z = –9 x + 2y + 2z = 5 x – 2y + 3z = –2 (1, 1, –10) (1, 2, –7) (–1, 2, 1) (–1, 1, 0)

OpenStudy (anonymous):

What is the equation of a vertical line passing through (3, –5)? y = 3 x = 3 y = –5 x = –5

OpenStudy (anonymous):

My guess is A, B, C

OpenStudy (anonymous):

Please someone help! My last exam questions and I want to go to sleep!!

OpenStudy (anonymous):

Just looking at your last question, that is incorrect. A vertical line must have as its equation x = b, where b is a real number. Why not plot the point and draw a vertical line through that point.

OpenStudy (anonymous):

Welp. I couldn't wait any longer. It doesn't say the answer but I know what I got right and wrong. It was A, not B, and not A